OpenStudy (anonymous):

Find the domain and range: f(x)=sqrt(32+4x^2) Calculate the exact value(s) of x when f(x)=6.

OpenStudy (anonymous):

that will give you your domain

OpenStudy (anonymous):

solve 32+4x^2>or equal to 0

OpenStudy (anonymous):

I gathered no solution.

OpenStudy (anonymous):

4x^2>=-32 x^2>=-8 are you sure it's no solutions?

OpenStudy (anonymous):

Isnt domain all real numbers?

OpenStudy (anonymous):

The domain is all real numbers. The range is f(x) is greater than sqrt(32)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=f%28x%29%3Dsqrt%2832%2B4x%5E2%29 4x^2=4 X^2=1 X=+-1

OpenStudy (anonymous):

forgot to put beginning sqrt(32+4x^2)=6 32+4x^2=36

OpenStudy (anonymous):

wrt domain you only have to find any values of x that would make 32+4x^2 negative.... and there are none! the only way that expression could even be 0 is if the squared term equals -32. no calculations necessary

OpenStudy (anonymous):

the system says incorrect...no luck with the domain. :\

OpenStudy (anonymous):

domain should be all real numbers because you can never have a zero under the square root bcz of the square on x

OpenStudy (anonymous):

also u cant have ive under the square root bcz of x^2

OpenStudy (anonymous):

-ive*